OpenStudy (anonymous):

Austin wants to sketch the graph of a circle represented by the given equation. (x + 3)2 + (y - 5)2 = 36 Using complete sentences, describe the steps used to sketch the circle on a coordinate grid.

Directrix (directrix):

(x + 3)^2 + (y - 5)^2 = 36 ------------------------- 1. Write the equation in the form (x - (-3))^2 + (y + (-5))^2 = 6^2 2. Obtain the center (-3, 5) and radius 6. 3. Mark the point (-3, 5) on the coordinate axis. 4. Calculate and plot the points around the center using given radius 6. 5. Draw a circle through the points plotted at distance 6 from the center.